# This is the extended version of the conference version at FoSSaCS 2022 [BFPZ22] [a] [a] [a]

### v1

@@ 5.3. Quality-optimal probability-raising causes
By definition of SPR causes any subset of $\mathcal{C}$ is a SPR cause and thus also $\mathsf{CanCause}$. Optimality is a consequence as $\mathsf{CanCause}$ even yields path-wise optimal coverage in the following sense. If $C$ is any SPR cause then $C\subseteq\mathcal{C}$ by definition and for each path $\pi$ in $\mathcal{M}$:
@@
We regard the given Markov chain $\mathcal{M}$ as an MDP with a singleton action set $\mathit{Act}=\{\alpha\}$. As $\mathcal{M}$ has SPR causes, the set $\mathcal{C}$ of states that constitute a singleton SPR cause is nonempty. We may assume that $\mathcal{M}$ has no non-trivial (i.e., cyclic) bottom strongly connected components as we may collapse them. Let $w_{c}$ == $\mathrm{Pr}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$. We switch from $\mathcal{M}$ to a new MDP $\mathcal{K}$ with state space $S_{\mathcal{K}}=S\cup\{\mathsf{eff}_{\mathsf{cov}},\mathsf{noeff}_{\mathsf{fp}}\}$ with fresh states $\mathsf{noeff}_{\mathsf{fp}}$ and $\mathsf{eff}_{\mathsf{cov}}$ and the action set $\mathit{Act}_{\mathcal{K}}=\{\alpha,\gamma\}$. The MDP $\mathcal{K}$ arises from $\mathcal{M}$ by adding (i) for each state $c\in\mathcal{C}$ a fresh state-action pair $(c,\gamma)$ with $P_{\mathcal{K}}(c,\gamma,\mathsf{eff}_{\mathsf{cov}})=w_{c}$ and $P_{\mathcal{K}}(c,\gamma,\mathsf{noeff}_{\mathsf{fp}})=1{-}w_{c}$ and (ii) reset transitions to $\mathsf{init}$ with action label $\alpha$ from the new state $\mathsf{noeff}_{\mathsf{fp}}$ and all terminal states of $\mathcal{M}$, i.e., $P_{\mathcal{K}}(\mathsf{noeff}_{\mathsf{fp}},\alpha,\mathsf{init})=1$ and $P_{\mathcal{K}}(s,\alpha,\mathsf{init})=1$ for $s\in\mathsf{Eff}$ or if $s$ is a terminal non-effect state of $\mathcal{M}$. So, exactly $\mathsf{eff}_{\mathsf{cov}}$ is terminal in $\mathcal{K}$, and $\mathit{Act}_{\mathcal{K}}(c)=\{\alpha,\gamma\}$ for $c\in\mathcal{C}$, while $\mathit{Act}_{\mathcal{K}}(s)=\{\alpha\}$ for all other states $s$. Intuitively, taking action $\gamma$ in state $c\in\mathcal{C}$ selects $c$ to be a cause state. The states in $\mathsf{Eff}$ represent uncovered effects in $\mathcal{K}$, while $\mathsf{eff}_{\mathsf{cov}}$ stands for covered effects.
@@
We define the subset $D\subseteq S_{\mathcal{N}}$ by
@@
Note that $\mathsf{init}_{\mathcal{N}}\in D$. For each $s\in D$, we further define
@@
Finally, let $E\subseteq D$ be the set of states that are reachable from $\mathsf{init}_{\mathcal{N}}$ when only choosing actions from $\mathit{Act}^{\min}(\cdot)$. Note that $E$ does not contain any states from $\mathcal{C}$.
@@
We now construct a further MDP $\mathcal{K}$. The idea is that $\mathcal{K}$ behaves like $\mathcal{N}$ after initially a scheduler is forced to choose a probability distribution over state-action pairs from $\Pi$. In this way, $\mathsf{Eff}$ is reached with positive probability under all schedulers. Given an SPR cause, we will observe that for the f-score of this cause under a scheduler, it is only important how large the probabilities with which state action pairs from $\Pi$ are chosen are relative to each other while the absolute values are not important. Due to this observation, for each SPR cause $C$ and for each scheduler $\mathfrak{S}$ for $\mathcal{N}$ that reaches $\mathsf{Eff}$ with positive probability, we can then construct a scheduler for $\mathcal{K}$ that leads to the same recall and precision of $C$.
@@
Proof of Claim 2. We first prove the direction $\Rightarrow$”. So, let $C$ be an SPR cause for $\mathsf{Eff}$ in $\mathcal{N}$ with $\mathit{fscore}(C)>\vartheta$. As first observation we have that in order to prove (17) for all schedulers $\mathfrak{T}$ for $\mathcal{K}$, it suffices to consider schedulers $\mathfrak{T}$ that start with a deterministic choice for state $\mathsf{init}_{\mathcal{K}}$ and then behave in an arbitrary way.
@@
Consider a scheduler $\mathfrak{T}$ for $\mathcal{K}$ with a deterministic choice $\mathfrak{T}(\mathsf{init}_{\mathcal{K}})(\beta_{(s,\alpha)})=1$ where $(s,\alpha)\in\Pi$. To construct an analogous scheduler $\mathfrak{S}$ of $\mathcal{N}$, we pick an MD-scheduler $\mathfrak{U}$ of the sub-MDP $\mathcal{N}^{\min}_{E}$ of $\mathcal{N}$ induced by the state-action pairs $(u,\beta)$ where $u\in E$ and $\beta\in\mathit{Act}^{\min}(u)$ such that there is a $\mathfrak{U}$-path from $\mathsf{init}_{\mathcal{N}}$ to state $s$.
@@
As $\mathfrak{S}$ and $\mathfrak{T}$ behave identically after choosing the state-action pair $(s,\alpha)\in\Pi$ or the corresponding action $\beta_{(s,\alpha)}$, respectively, which implies that
@@
As $\mathfrak{S}$ leaves the sub-MDP $\mathcal{N}^{\min}_{E}$ with probability $\lambda>0$, we have $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond\mathsf{Eff})>0$. By Lemma 40, we can conclude that
$2(1{-}\vartheta)\mathsf{tp}^{\mathfrak{S}}_{\mathcal{N}}-\vartheta\mathsf{fp}^{\mathfrak{S}}_{\mathcal{N}}-\vartheta\mathsf{fn}^{\mathfrak{S}}_{\mathcal{N}}>0$
@@
$2(1{-}\vartheta)\mathsf{tp}^{\mathfrak{T}}_{\mathcal{K}}-\vartheta\mathsf{fp}^{\mathfrak{T}}_{\mathcal{K}}-\vartheta\mathsf{fn}^{\mathfrak{T}}_{\mathcal{K}}>0$
@@
$2(1{-}\vartheta)\mathsf{tp}^{\mathfrak{T}}_{\mathcal{K}}-\vartheta\mathsf{fp}^{\mathfrak{T}}_{\mathcal{K}}-\vartheta\mathsf{fn}^{\mathfrak{T}}_{\mathcal{K}}>0$
@@
We now construct a stochastic shortest path game (see [PB99]) to check whether there is a subset $C\subseteq\mathcal{C}$ in $\mathcal{K}$ such that (17) holds. Such a game is played on an MDP-like structure with the only difference that the set of states is partitioned into two sets indicating which player controls which states.
The game $\mathcal{G}$ has states $(S_{\mathcal{K}}\times\{\mathit{yes},\mathit{no}\})\cup\mathcal{C}\times\{\mathit{choice}\}$. On the subset $S_{\mathcal{K}}\times\{\mathit{yes}\}$, all available actions and transition probabilities are just as in $\mathcal{K}$ and this copy of $\mathcal{K}$ cannot be left. More formally, for all $s,t\in S_{\mathcal{K}}$ and $\alpha\in\mathit{Act}_{\mathcal{K}}(s)$, we have $\mathit{Act}_{\mathcal{G}}((s,\mathit{yes}))=\mathit{Act}_{\mathcal{K}}(s)$ and $P_{\mathcal{G}}((s,\mathit{yes}),\alpha,(t,\mathit{yes}))=P_{\mathcal{K}}(s,\alpha,t)$.
In the $\mathit{no}$”-copy, the game also behaves like $\mathcal{G}$ but when a state in $\mathcal{C}$ would be entered, the game moves to a state in $\mathcal{C}\times\{\mathit{choice}\}$ instead. In a state of the form $(c,\mathit{choice})$ with $c\in\mathcal{C}$, two actions $\alpha$ and $\beta$ are available. Choosing $\alpha$ leads to the state $(c,\mathit{yes})$ while choosing $\beta$ leads to $(c,\mathit{no})$ with probability $1$.
@@
$\mathrm{NP}$-membership for Markov chains.
$\mathrm{NP}$-membership for all three problems within Markov chains is straightforward as we may non-deterministically guess a cause and check in polynomial time whether it constitutes a GPR cause and satisfies the threshold condition for the recall, coverage ratio or f-score.
$\mathrm{NP}$-hardness of GPR-recall and GPR-covratio.
With arguments as in the proof of Lemma 35, the problems GPR-recall and GPR-covratio are polynomially interreducible for Markov chains. Thus, it suffices to prove NP-hardness of GPR-recall. For this, we provide a polynomial reduction from the knapsack problem. The input of the latter are sequences $A_{1},\ldots,A_{n},A$ and $B_{1},\ldots,B_{n},B$ of positive natural numbers and the task is to decide whether there exists a subset $I$ of $\{1,\ldots,n\}$ such that
@@
$\mathrm{NP}$-hardness of GPR-f-score.
Using similar ideas, we also provide a polynomial reduction from the knapsack problem. Let $A,A_{1},\ldots,A_{n},B,B_{1},\ldots,B_{n}$ be an input for the knapsack problem. We replace the $A$-sequence with $a,a_{1},\ldots,a_{n}$ where $a=\frac{A}{N}$ and $a_{i}=\frac{A_{i}}{N}$ where $N$ is as before. The topological structure of the Markov chain that we are going to construct is the same as in the NP-hardness proof for GPR-recall.
@@
As we can decide for a given cause candidate $\mathsf{Cause}$ whether it is a SPR cause in $\mathrm{P}$ or a GPR cause in $\mathrm{coNP}$, this also yields an algorithm for finding optimal causes for $f$. Given an MDP $\mathcal{M}$ with terminal effect set $\mathsf{Eff}$ and quality measure $f$ as algebraic function we consider each cause candidate $\mathsf{Cause}$, check whether it is a PR cause (SPR or GPR) and consider the decision problem $f(\mathsf{Cause})\leq\vartheta$. As all of these steps have a complexity upper bound of $\mathrm{PSPACE}$ and we only need to save the best cause candidate so far with its value $f(\mathsf{Cause})$, this results in polynomial space as well.
@@ 6.1. Sets of states as causes
There is one small caveat that we want to mention here: If the effect $\mathsf{rEff}$ is a reachability property $\lozenge\mathsf{Eff}$ for a set of states $\mathsf{Eff}\subseteq S$, then this new definition allows that a GPR/SPR cause $\mathsf{Cause}$ is not disjoint from the set of states $\mathsf{Eff}$. If two sets $\mathsf{Cause},\mathsf{Eff}\subseteq S$ are disjoint, however, then $\mathsf{Cause}$ is a GPR/SPR cause for $\mathsf{Eff}$ according to Definition 3 iff $\mathsf{Cause}$ is a reachability GPR/SPR cause for the $\omega$-regular event $\lozenge\mathsf{Eff}$ according to the new definition. As we now view the effect as the $\omega$-regular property on infinite executions, one can, nevertheless, argue that the temporal priority (C2) is captured by the new definition because the cause will be reached after finitely many steps if it is reached. We will address problems with this interpretation and a stronger notion of temporal priority in Section 6.1.3.

### v2

@@ 5.3. Quality-optimal probability-raising causes
By definition of SPR causes any subset $C\subseteq\mathcal{C}$ satisfying $\mathrm{Pr}^{\max}(\neg C\,\mathrm{U}\,c)$ for each $c\in C$ constitutes an SPR cause and thus $\mathsf{CanCause}$ is also an SPR cause. Optimality is a consequence as $\mathsf{CanCause}$ even yields path-wise optimal coverage in the following sense. If $C$ is any SPR cause then $C\subseteq\mathcal{C}$ by definition and for each path $\pi$ in $\mathcal{M}$:
@@
We regard the given Markov chain $\mathcal{M}$ as an MDP with a singleton action set $\mathit{Act}=\{\alpha\}$. As $\mathcal{M}$ has SPR causes, the set $\mathcal{C}$ of states that constitute a singleton SPR cause is nonempty. We may assume that $\mathcal{M}$ has no non-trivial (i.e., cyclic) bottom strongly connected components as we may collapse them. Let $w_{c}$ == $\mathrm{Pr}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$.
We switch from $\mathcal{M}$ to a new MDP $\mathcal{K}$ with state space $S_{\mathcal{K}}=S\cup\{\mathsf{eff}_{\mathsf{cov}},\mathsf{noeff}_{\mathsf{fp}}\}$ with fresh states $\mathsf{noeff}_{\mathsf{fp}}$ and $\mathsf{eff}_{\mathsf{cov}}$ and the action set $\mathit{Act}_{\mathcal{K}}=\{\alpha,\gamma\}$. The MDP $\mathcal{K}$ arises from $\mathcal{M}$ by adding
(i):
for each SPR state $c\in\mathcal{C}$ a fresh state-action pair $(c,\gamma)$ such that $P_{\mathcal{K}}(c,\gamma,\mathsf{eff}_{\mathsf{cov}})=w_{c}$ and $P_{\mathcal{K}}(c,\gamma,\mathsf{noeff}_{\mathsf{fp}})=1{-}w_{c}$ and
(ii):
reset transitions to $\mathsf{init}$ with action label $\alpha$ from the new state $\mathsf{noeff}_{\mathsf{fp}}$ and all terminal states of $\mathcal{M}$, i.e., $P_{\mathcal{K}}(\mathsf{noeff}_{\mathsf{fp}},\alpha,\mathsf{init})=1$ and $P_{\mathcal{K}}(s,\alpha,\mathsf{init})=1$ for $s\in\mathsf{Eff}$ or if $s$ is a terminal non-effect state of $\mathcal{M}$.
So, exactly $\mathsf{eff}_{\mathsf{cov}}$ is terminal in $\mathcal{K}$, and $\mathit{Act}_{\mathcal{K}}(c)=\{\alpha,\gamma\}$ for $c\in\mathcal{C}$, while $\mathit{Act}_{\mathcal{K}}(s)=\{\alpha\}$ for all other states $s$. Intuitively, taking action $\gamma$ in state $c\in\mathcal{C}$ selects $c$ to be a cause state. The states in $\mathsf{Eff}$ represent uncovered effects in $\mathcal{K}$, while $\mathsf{eff}_{\mathsf{cov}}$ stands for covered effects.
@@
We define the subset of states from which $\mathsf{Eff}$ can be avoided as $D\subseteq S_{\mathcal{N}}$ by
@@
Note that $\mathsf{init}_{\mathcal{N}}\in D$. For each $s\in D$, we further define the set of actions minimizing the reachability probability of $\mathsf{Eff}$ from $s$ by
@@
Finally, let $E\subseteq D$ be the set of states that are reachable from $\mathsf{init}_{\mathcal{N}}$ when only choosing actions from $\mathit{Act}^{\min}(\cdot)$. Note that $E$ does not contain any states from $\mathcal{C}$, meaning no state in $E$ constitutes a singleton SPR cause.
@@
We now construct a further MDP $\mathcal{K}$. The idea is that $\mathcal{K}$ behaves like the end-component free MDP $\mathcal{N}$ after initially a scheduler is forced to choose a probability distribution over state-action pairs from $\Pi$. In this way, $\mathsf{Eff}$ is reached with positive probability under all schedulers. Given an SPR cause, we will observe that for the f-score of this cause under a scheduler, it is only important how large the probabilities with which state action pairs from $\Pi$ are chosen are relative to each other while the absolute values are not important. Due to this observation, for each SPR cause $C$ and for each scheduler $\mathfrak{S}$ for $\mathcal{N}$ that reaches $\mathsf{Eff}$ with positive probability, we can then construct a scheduler for $\mathcal{K}$ that leads to the same recall and precision of $C$.
@@
Proof of Claim 2. We first prove the direction $\Rightarrow$”. So, let $C$ be an SPR cause for $\mathsf{Eff}$ in end-component free MDP $\mathcal{N}$ with $\mathit{fscore}(C)>\vartheta$. As first observation we have that in order to prove (17) for all schedulers $\mathfrak{T}$ for $\mathcal{K}$, it suffices to consider schedulers $\mathfrak{T}$ that start with a deterministic choice for state $\mathsf{init}_{\mathcal{K}}$ and then behave in an arbitrary way.
@@
Recall the set $\mathit{Act}^{\min}(s)$ of actions minimizing the reachability probability of $\mathsf{Eff}$ from $s$. Consider a scheduler $\mathfrak{T}$ for $\mathcal{K}$ with a deterministic choice $\mathfrak{T}(\mathsf{init}_{\mathcal{K}})(\beta_{(s,\alpha)})=1$ where $(s,\alpha)\in\Pi$. To construct an analogous scheduler $\mathfrak{S}$ of $\mathcal{N}$, we pick an MD-scheduler $\mathfrak{U}$ of the sub-MDP $\mathcal{N}^{\min}_{E}$ of $\mathcal{N}$ induced by the state-action pairs $(u,\beta)$ where $u\in E$ and $\beta\in\mathit{Act}^{\min}(u)$ such that there is a $\mathfrak{U}$-path from $\mathsf{init}_{\mathcal{N}}$ to state $s$.
@@
$\mathfrak{S}$ and $\mathfrak{T}$ behave identically after choosing the state-action pair $(s,\alpha)\in\Pi$ or the corresponding action $\beta_{(s,\alpha)}$, respectively, which implies that
@@
Recall the sub-MDP $\mathcal{N}^{\min}_{E}$ consisting of $E$ and the actions in $\mathit{Act}^{\min}(\cdot)$. As $\mathfrak{S}$ leaves the sub-MDP $\mathcal{N}^{\min}_{E}$ with probability $\lambda>0$, we have $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond\mathsf{Eff})>0$. By Lemma 40, we can conclude that
$2(1{-}\vartheta)\mathsf{tp}^{\mathfrak{S}}_{\mathcal{N}}-\vartheta\mathsf{fp}^{\mathfrak{S}}_{\mathcal{N}}-\vartheta\mathsf{fn}^{\mathfrak{S}}_{\mathcal{N}}>0.$
@@
$2(1{-}\vartheta)\mathsf{tp}^{\mathfrak{T}}_{\mathcal{K}}-\vartheta\mathsf{fp}^{\mathfrak{T}}_{\mathcal{K}}-\vartheta\mathsf{fn}^{\mathfrak{T}}_{\mathcal{K}}>0.$
@@
$2(1{-}\vartheta)\mathsf{tp}^{\mathfrak{T}}_{\mathcal{K}}-\vartheta\mathsf{fp}^{\mathfrak{T}}_{\mathcal{K}}-\vartheta\mathsf{fn}^{\mathfrak{T}}_{\mathcal{K}}>0.$
@@
Recall the set of singleton SPR causes $\mathcal{C}$. We now construct a stochastic shortest path game (see [PB99]) to check whether there is a subset $C\subseteq\mathcal{C}$ such that (17) holds in the EC-free MDP $\mathcal{K}$ in which visiting effect states always has a non-zero probability. Such a game is played on an MDP-like structure with the only difference that the set of states is partitioned into two sets indicating which player controls which states.
The game $\mathcal{G}$ has states $(S_{\mathcal{K}}\times\{\mathit{yes},\mathit{no}\})\cup\mathcal{C}\times\{\mathit{choice}\}$. On the subset $S_{\mathcal{K}}\times\{\mathit{yes}\}$, all actions and transition probabilities are just as in $\mathcal{K}$ and this copy of $\mathcal{K}$ cannot be left. Formally, for all $s,t\in S_{\mathcal{K}}$ and $\alpha\in\mathit{Act}_{\mathcal{K}}(s)$, we have $\mathit{Act}_{\mathcal{G}}((s,\mathit{yes}))=\mathit{Act}_{\mathcal{K}}(s)$ and $P_{\mathcal{G}}((s,\mathit{yes}),\alpha,(t,\mathit{yes}))=P_{\mathcal{K}}(s,\alpha,t)$.
In the $\mathit{no}$”-copy, the game also behaves like $\mathcal{K}$ but when a state in $\mathcal{C}$ would be entered, the game moves to a state in $\mathcal{C}\times\{\mathit{choice}\}$ instead. In a state of the form $(c,\mathit{choice})$ with $c\in\mathcal{C}$, two actions $\alpha$ and $\beta$ are available. Choosing $\alpha$ leads to the state $(c,\mathit{yes})$ while choosing $\beta$ leads to $(c,\mathit{no})$ with probability $1$.
@@
$\mathrm{NP}$-membership for Markov chains. $\mathrm{NP}$-membership for all three problems within Markov chains is straightforward as we may non-deterministically guess a cause and check in polynomial time whether it constitutes a GPR cause and satisfies the threshold condition for the recall, coverage ratio or f-score.
$\mathrm{NP}$-hardness of GPR-recall and GPR-covratio. With arguments as in the proof of Lemma 35, the problems GPR-recall and GPR-covratio are polynomially interreducible for Markov chains. Thus, it suffices to prove NP-hardness of GPR-recall. For this, we provide a polynomial reduction from the knapsack problem. The input of the latter are sequences $A_{1},\ldots,A_{n},A$ and $B_{1},\ldots,B_{n},B$ of positive natural numbers and the task is to decide whether there exists a subset $I$ of $\{1,\ldots,n\}$ such that
@@
$\mathrm{NP}$-hardness of GPR-f-score. Using similar ideas, we also provide a polynomial reduction from the knapsack problem. Let $A,A_{1},\ldots,A_{n},B,B_{1},\ldots,B_{n}$ be an input for the knapsack problem. We replace the $A$-sequence with $a,a_{1},\ldots,a_{n}$ where $a=\frac{A}{N}$ and $a_{i}=\frac{A_{i}}{N}$ where $N$ is as before. The topological structure of the Markov chain that we are going to construct is the same as in the NP-hardness proof for GPR-recall.
@@
As we can decide for a given cause candidate $\mathsf{Cause}$ whether it is a SPR cause in $\mathrm{P}$ or a GPR cause in $\mathrm{coNP}$, this also yields an algorithm for finding optimal causes for $f$. Given an MDP $\mathcal{M}$ with terminal effect set $\mathsf{Eff}$ and quality measure $f$ as an algebraic function we consider each cause candidate $\mathsf{Cause}$, check whether it is a PR cause (SPR or GPR) and consider the decision problem $f(\mathsf{Cause})\leq\vartheta$. As all of these steps have a complexity upper bound of $\mathrm{PSPACE}$ and we only need to save the best cause candidate so far with its value $f(\mathsf{Cause})$, this results in polynomial space as well.
@@ 6.1. Sets of states as causes
There is one small caveat that we want to mention here: If the effect $\mathsf{rEff}$ is a reachability property $\lozenge\mathsf{Eff}$ for a set of states $\mathsf{Eff}\subseteq S$, then this new definition allows for GPR/SPR causes $\mathsf{Cause}$ not disjoint from the set of states $\mathsf{Eff}$. If two sets $\mathsf{Cause},\mathsf{Eff}\subseteq S$ are disjoint, however, then $\mathsf{Cause}$ is a GPR/SPR cause for $\mathsf{Eff}$ according to Definition 3 iff $\mathsf{Cause}$ is a reachability GPR/SPR cause for the $\omega$-regular event $\lozenge\mathsf{Eff}$ according to the new definition. As we now view the effect as the $\omega$-regular property on infinite executions, one can, nevertheless, argue that the temporal priority (C2) is captured by the new definition since the cause will be reached after finitely many steps if it is reached. We will address problems with this interpretation and a stronger notion of temporal priority in Section 6.1.3.